OpenStudy (anonymous):

4x+11=7 Solve the value of -8x. Show your work

OpenStudy (anonymous):

@zaynahf here

OpenStudy (anonymous):

If you have: 4x+11=7 You need to isolate the x term. How can you start out with this?

OpenStudy (anonymous):

i didnt get the -8x i dont know if i should sub x for -8

OpenStudy (anonymous):

as far as i know, you have to find out what x is first, do you know the value of x?

OpenStudy (anonymous):

4x+11=7 Solve the value of -8x. Show your work does that mean x is -8

OpenStudy (anonymous):

No, like i said before, find what x is from 4x+11=7.

OpenStudy (anonymous):

add 4 and 11 =15

OpenStudy (anonymous):

No, you dont add 4. Subtract 11 from both sides

OpenStudy (anonymous):

x+11=3

OpenStudy (anonymous):

If you subtract 11: 4x+11=7 -11 -11 ------------ 4x=-4 See how i got that?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Ok now divide by 4 on bothsides

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

do u do online school?

OpenStudy (anonymous):

??? is that the anwser

OpenStudy (anonymous):

Yeah, thats right. No, i dont do online school. Now plug in -1 for x in the -8x

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

Thats not right.. remember that 2 negatives make a positive

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

Yeah, thats the answer

OpenStudy (anonymous):

i solved an equation how to i graph it 6x>12x

OpenStudy (anonymous):

thats a new question

OpenStudy (anonymous):

Please close this, and open a new question/..